# Corresponding Authors

### v1

@@ Abstract
Video frame interpolation (VFI) aims to synthesize intermediate frames in between existing frames to enhance visual smoothness and quality. Beyond the conventional methods based on the reconstruction loss, recent works employ the high quality generative models for perceptual quality. However, they require complex training and large computational cost for modeling on the pixel space. In this paper, we introduce disentangled Motion Modeling (MoMo), a diffusion-based approach for VFI that enhances visual quality by focusing on intermediate motion modeling. We propose disentangled two-stage training process, initially training a frame synthesis model to generate frames from input pairs and their optical flows. Subsequently, we propose a motion diffusion model, equipped with our novel diffusion U-Net architecture designed for optical flow, to produce bi-directional flows between frames. This method, by leveraging the simpler low-frequency representation of motions, achieves superior perceptual quality with reduced computational demands compared to generative modeling methods on the pixel space. Our method surpasses state-of-the-art methods in perceptual metrics across various benchmarks, demonstrating its efficacy and efficiency in VFI. Our code is available at: https://github.com/JHLew/MoMo
Figure 1: Video Frame Interpolation result of our proposed method (MoMo) with comparison to state-of-the-art methods. Our approach produces the most visually pleasant result, owing to proper modeling of the intermediate motion.
@@ 1 Introduction
Video Frame Interpolation (VFI) is a crucial task in computer vision that aims to synthesize absent frames between existing ones in a video. It has a wide spectrum of applications, such as generating slow motion [23], compressing video [68], and producing animation [63]. Its ultimate goal is to elevate the visual quality of videos through enhanced motion smoothness and image sharpness. Motions, represented by optical flows [65, 69, 66, 21] and realized by warping, have been central to VFI’s development as the majority of recent innovations in VFI have been accomplished along with advances in intermediate motion estimation [20, 48, 49, 32, 61, 24, 71, 22, 70, 28, 52, 72, 2, 25, 47, 5].
However, these approaches often result in perceptually unsatisfying outcomes due to their reliance on $L_{1}$ or $L_{2}$ objectives, leading to high PSNR scores but poor perceptual quality [7, 30, 73]. To address this matter, recent advancements [6, 23, 45, 4] have explored the use of deep feature spaces to achieve improved quality in terms of human perception [26, 73]. Additionally, the integration of generative models into VFI [31, 8, 67] has introduced novel pathways for improving the visual quality of videos but has primarily focused on modeling pixels or latent spaces directly, which demands high computational resources.
We introduce disentangled Motion Modeling (MoMo), a perception-oriented approach for VFI, focusing on the modeling of intermediate motions rather than direct pixel generation. We employ a diffusion model [18] to generate bi-directional optical flow maps, marking the first use of generative modeling for motion in VFI. We propose a disentangled two-stage training process that includes the first stage of training a frame synthesis model and fine-tuning of an optical flow estimator [66]. This flow estimator then provides ground-truths in the second stage of training, of our motion diffusion model, which, during inference, generates the motions needed for the synthesis network to create the final interpolated frame. We propose a novel architecture for our motion diffusion model, considering the nature of optical flows, enhancing both computational efficiency and performance.
Our experiments validate the effectiveness and efficiency of our proposed architecture and training scheme, demonstrating superior performance across various benchmarks in terms of perceptual metrics. By focusing on the generative modeling of motions rather than the direct generation of frames, our approach achieves improvements in visual quality, addressing the core objective of VFI.
@@
  • We propose a two-stage process, disentangling frame synthesis and motion modeling, the two crucial components in VFI.
  • We introduce a novel diffusion-model architecture suitable for optical flow modeling, boosting efficiency and quality.
@@ 2.1 Flow-based Video Frame Interpolation
In deep learning-based Video Frame Interpolation (VFI), optical flow-based methods have recently become prominent, and these methods typically follow a common two-stage process. Initially, they estimate the flows needed to create the target intermediate frame, which involves warping of the input frames with these flows. Then, a synthesis network merges the warped images to produce the final frame. The direction of the required flows varies based on the warping technique, either towards or from the target frame. Forward warping, utilized in various studies [45, 44, 25, 24, 20, 43], exploit flows directed towards the target frame. The flows are achieved by linearly adjusting the bi-directional flows between input frames according to the target frame’s intermediate timestep. Conversely, backward warping-based methods focus on estimating flows originating from the target frame, with early approaches also relying on bi-directional flows between input frames [23, 2, 70, 61], which are often converted with flow reversal [70].
Recent advancements in VFI have focused on directly predicting intermediate flows from input frames [22, 28, 35, 72, 32, 48, 49, 47], sparking specialized architectures such as bilateral cost volumes [48, 49, 47] to improve flow accuracy. As VFI quality progresses with flow estimation enhancements, our work aims to improve intermediate flow predictions further. Unlike most methods that rely on reconstruction loss for end-to-end training and some that add flow distillation loss [22, 28, 35, 32] as an auxiliary loss, our approach employs disentangled and direct supervision solely for flow estimation, marking an innovation in VFI research.
@@ 2.2 Perception-oriented Restoration
Conventional restoration methods in computer vision, including VFI, focused on minimizing $L_{1}$ or $L_{2}$ distances, often result in blurry images [57, 30, 58, 39] due to prioritizing pixel accuracy over human visual perception. Recent studies have shifted towards deep feature spaces for reconstruction loss [26] and evaluation metrics [73, 10], demonstrating that these align better with human judgment. These approaches emphasize perceptual quality over traditional metrics like PSNR, signaling a move towards more visually appealing, photo-realistic image synthesis.
Moreover, generative models have also been acknowledged for their ability to enhance visual quality in restoration tasks. Early work by Ledig et al. [30] utilized generative adversarial networks (GANs) [16] to produce visually superior restoration results, combining the adversarial loss with traditional reconstruction losses. Beyond GANs, other generative models such as normalizing flows [53, 11, 12] and diffusion models [64, 18] have demonstrated similar observations in improving perceptual quality for restoration tasks [36, 57]. In video frame interpolation (VFI), the adoption of generative models has also been explored [31, 67, 8]. LDMVFI [8], closely related to our work, use latent diffusion models [54] to enhance perceptual quality. Our method aligns with such innovations but uniquely focuses on generating optical flow maps, differing from prior generative approaches that directly model the pixels of the target frame.
@@ 2.3 Diffusion Models
Diffusion models [64, 18, 54] are popular generative models that consist of forward and reverse process. Initially, the forward process incrementally adds noise to the data $\textbf{x}_{0}$ over $T$ steps via a predefined Markov chain, resulting in $\textbf{x}_{T}$ that approximates a Gaussian noise. The diffused data $\textbf{x}_{t}$ is obtained through forward process:
@@
While a commonly used approach is to predict the noise as above ($\epsilon$-prediction), there are some alternative approachs, such as $\textbf{x}_{0}$-prediction [51] which predicts the data $\textbf{x}_{0}$ itself or v-prediction [59], which is beneficial for numerical stability.
Diffusion model synthesizes the data in an iterative manner following the backward process, resulting in high perceptual quality of image samples or video samples [9, 56, 19, 17]. Further, we are motivated by optical flow modeling with diffusion models in other tasks [60, 42], and aim to leverage the benefit of diffusion models for optical flow synthesis in video frame interpolation. Although existing work employs diffusion model for video frame interpolation task [8], our method synthesizes the intermediate optical flow rather than directly synthesizing the RGB frames.
@@ 3.1 Overview
Figure 2: Overview of our entire framework, MoMo. Our framework operates in two stages. Initially, we train a frame synthesis network and an optical flow model, with the latter providing ground-truth labels for the second stage. In the second stage of training, we focus on training a Motion Diffusion Model to predict bi-directional flow between frames. During inference, this model generates flow fields from the input frame pair, enabling the frame synthesis model to output the final target frame.
In this paper, we focus on Video Frame Interpolation (VFI) with the goal of synthesizing an intermediate frame $I_{\tau}$ between consecutive frames $I_{0}$ and $I_{1}$, where $10<\tau<$ 1. Our method MoMo adopts a two-stage training scheme to disentangle the training of motion modeling and frame synthesis (Fig. 2). In the first stage, we train a frame synthesis network to synthesize an RGB frame from neighboring frames and their bi-directional flows. Then, we fine-tune the optical flow estimator to enhance flow quality. In the second stage, the fine-tuned flow estimator serves as a teacher for training the motion diffusion model. During inference, this model generates intermediate motion, bi-directional flow maps in specific, which the synthesis network uses to produce the final RGB frame.
3.2 Synthesis Network and Teacher Flow Estimator
We propose a synthesis network $\mathcal{S}$, designed to accurately generate an intermediate target frame using a pair of input frames and their corresponding optical flows from the target frames. Specifically, given a frame pair of $I_{0},I_{1}$, and the target intermediate frame $I_{\tau}$, we first use an optical flow network $\mathcal{F}$ to obtain the bi-directional flow from the target frame to the input frames:
@@ 3.2 Synthesis Network and Teacher Flow Estimator
where $i$ denote the index of input frames. With the estimated flows and their corresponding frames from the target frame, we synthesize $\hat{I}_{\tau}$, which aims to recover the target frame $I_{\tau}$.
@@
We adopt pre-trained RAFT [66] for optical flow model $\mathcal{F}$, and train the synthesis network $\mathcal{S}$ from scratch. We choose an alternating optimization of two models, $\mathcal{S}$ and $\mathcal{F}$. We first fix $\mathcal{F}$ to the pre-trained state, and train the synthesis network $\mathcal{S}$. Once the training of $\mathcal{S}$ converges, we freeze it, and fine-tune $\mathcal{F}$. We tune the optical flow model $\mathcal{F}$ so that it could provide better estimations as the teacher in the next stage of training. Note that the flow estimator $\mathcal{F}$ is not used during inference, but serves its purpose as the teacher for intermediate motion modeling described in Sec. 3.3.
@@
For optimization, we compute loss on the final synthesized output $\hat{I}_{\tau}$, with a combination of three terms. First, we use the pixel reconstruction loss denoted as $\mathcal{L}_{1}$, to minimize the pixel-wise error between the synthesized frame and the target frame: $\mathcal{L}_{1}=||I_{\tau}-\hat{I}_{\tau}||_{1}$. As our ultimate goal is to produce frames of high-quality, we additionally use two losses that are known to be advantageous in promoting perceptual quality, along with $\mathcal{L}_{1}$. Following recent efforts [8, 4], we adopt the LPIPS-based perceptual reconstruction loss [73], denoted as $\mathcal{L}_{p}$, and also exploit the style loss [15] $\mathcal{L}_{G}$, as its effectiveness has been proved in a recent work [52]. By combining the three loss terms, we obtain $\mathcal{L}_{s}$, our perception-oriented reconstruction loss for high quality synthesis:
@@
Further details on our perception-oriented loss can be found in the Appendix.
3.2.1 Recurrent Synthesis
We build our synthesis network $\mathcal{S}$ to be of recurrent structure, motivated by the recent trend in video frame interpolation [61, 24, 52], due to its great efficiency. Let the number of recurrent process be $L-1$, and $\mathcal{S}$ can be expressed as recurrent application of a synthesis process $\mathcal{P}$:
$\mathcal{S}(I_{in},F_{\tau})=\mathcal{P}^{0}(\mathcal{P}^{1}(\cdots\mathcal{P}^{L-1}(\hat{I}^{L}_{\tau},I^{L-1}_{in},F^{L-1}_{\tau})\cdots,I^{1}_{in},F^{1}_{\tau}),I^{0}_{in},F^{0}_{\tau}),$ (6)
Next, we take frame $\hat{I}^{l+1}_{\tau}$ and use bicubic upsampling to match the size of level $l$, denoted as $\hat{I}^{l+1\rightarrow l}_{\tau}$. The upsampled frame $\hat{I}^{l+1\rightarrow l}_{\tau}$ and the two warped frames $I^{l}_{\tau\leftarrow 0},I^{l}_{\tau\leftarrow 1}$ are given to the synthesis module $\mathcal{G}$, which outputs a 4 channel output — 1 channel occlusion mask $M_{0}$ to blend $I_{\tau\leftarrow 0}$ and $I_{\tau\leftarrow 1}$, and 3 channel residual RGB values $\Delta\hat{I_{\tau}}$:
$M^{l}_{0},\Delta\hat{I}^{l}_{\tau}=\mathcal{G}(\hat{I}^{l+1\rightarrow l}_{\tau},I^{l}_{\tau\leftarrow 0},I^{l}_{\tau\leftarrow 1}).$ (8)
$\mathcal{P}^{l}(\hat{I}^{l+1}_{\tau},I_{in}^{l},F_{\tau}^{l})=\hat{I}^{l}_{\tau}=I^{l}_{\tau\leftarrow 0}\odot M^{l}_{0}+I^{l}_{\tau\leftarrow 1}\odot(1-M^{l}_{0})+\Delta\hat{I}^{l}_{\tau}.$ (9)
Note that $\hat{I}^{L}_{\tau}$ is not available for level $\mathcal{P}^{L-1}$, since it is of the highest level. Therefore we equally blend the two warped frames at level $l=L-1$ as a starting point: $\hat{I}^{L\rightarrow L-1}_{\tau}=I^{L-1}_{\tau\leftarrow 0}\odot 0.5+I^{L-1}_{\tau\leftarrow 1}\odot 0.5$. For our synthesis module $\mathcal{G}$, we use a simple 3-level hierarchy U-Net [55].

### v2

# Disentangled Motion Modeling for Video Frame Interpolation

@@ Abstract
Video Frame Interpolation (VFI) aims to synthesize intermediate frames between existing frames to enhance visual smoothness and quality. Beyond the conventional methods based on the reconstruction loss, recent works have employed generative models for improved perceptual quality. However, they require complex training and large computational costs for pixel space modeling. In this paper, we introduce disentangled Motion Modeling (MoMo), a diffusion-based approach for VFI that enhances visual quality by focusing on intermediate motion modeling. We propose a disentangled two-stage training process. In the initial stage, frame synthesis and flow models are trained to generate accurate frames and flows optimal for synthesis. In the subsequent stage, we introduce a motion diffusion model, which incorporates our novel U-Net architecture specifically designed for optical flow, to generate bi-directional flows between frames. By learning the simpler low-frequency representation of motions, MoMo achieves superior perceptual quality with reduced computational demands compared to the generative modeling methods on the pixel space. MoMo surpasses state-of-the-art methods in perceptual metrics across various benchmarks, demonstrating its efficacy and efficiency in VFI.
Figure 1: Video frame interpolation results of our proposed method called MoMo with comparison to state-of-the-art methods. MoMo produces the most visually pleasant result, owing to proper modeling of the intermediate motion.
Corresponding authorsCode available at: https://github.com/JHLew/MoMo
@@ 1 Introduction
Video Frame Interpolation (VFI) is a crucial task in computer vision that aims to synthesize absent frames between existing ones in a video. It has a wide spectrum of applications, such as slow motion generation (Jiang et al. 2018), video compression (Wu, Singhal, and Krahenbuhl 2018), and animation production (Siyao et al. 2021). Its ultimate goal is to elevate the visual quality of videos through enhanced motion smoothness and image sharpness. Motions, represented by optical flows (Sun et al. 2018; Teed and Deng 2020) and realized by warping, have been central to VFI’s development as recent innovations in VFI have mostly been accomplished along with advances in intermediate motion estimation (Xu et al. 2019; Chi et al. 2020; Park et al. 2020; Park, Lee, and Kim 2021; Sim, Oh, and Kim 2021; Reda et al. 2022; Jin et al. 2023).
However, these approaches often result in perceptually unsatisfying outcomes due to their reliance on $L_{1}$ or $L_{2}$ objectives, leading to high PSNR scores yet poor perceptual quality (Ledig et al. 2017; Zhang et al. 2018). To address this matter, recent advancements (Choi et al. 2020; Jiang et al. 2018; Niklaus and Liu 2020; Chen and Zwicker 2022) have explored the use of deep feature spaces to achieve improved quality in terms of human perception (Johnson, Alahi, and Fei-Fei 2016; Zhang et al. 2018). Additionally, the integration of generative models into VFI (Danier, Zhang, and Bull 2024; Wu et al. 2024) has introduced novel pathways for improving the visual quality of videos but has primarily focused on modeling pixels or latent spaces directly, which demands high computational resources.
We introduce disentangled Motion Modeling (MoMo), a perception-oriented approach for VFI, focusing on the modeling of intermediate motions rather than direct pixel generation. Here, we employ a diffusion model (Ho, Jain, and Abbeel 2020) to generate bi-directional optical flow maps, marking the first use of generative modeling for motion in VFI. We propose to disentangle the training of frame synthesis and intermediate motion prediction into a two-stage process: the initial stage includes the training of a frame synthesis model and fine-tuning of an optical flow model (Teed and Deng 2020). The frame synthesis model is designed to correctly synthesize an RGB frame given a pair of frames and their corresponding flow maps. In the subsequent stage of training, we train our motion diffusion model, which generates the intermediate motions the frame synthesis model uses to create the final interpolated frame during inference. In this stage of training, the optical flow model fine-tuned in the first stage serves as a teacher to provide pseudo-labels for the motion diffusion model. We also propose a novel architecture for our motion diffusion model, inspired by the nature of optical flows, enhancing both computational efficiency and performance.
Our experiments validate the effectiveness and efficiency of our proposed training scheme and architecture, demonstrating superior performance across various benchmarks in terms of perceptual metrics, with approximately 70×70$\times$70 × faster runtime compared to the existing diffusion-based VFI method (Danier, Zhang, and Bull 2024). By prioritizing the generative modeling of motions, our approach enhances visual quality, effectively addressing the core objective of VFI.
@@
  • We propose to disentangle the training of frame synthesis and intermediate motion modeling into a two-stage process, which are the crucial components in VFI.
  • We introduce a novel diffusion model architecture suitable for optical flow modeling, boosting efficiency and quality.
@@ 2.1 Flow-based Video Frame Interpolation
In deep learning-based Video Frame Interpolation (VFI), optical flow-based methods have recently become prominent, typically following a common two-stage process. First, the flows to or from the target intermediate frame is estimated, which involves warping of the input frame pair with the estimated flows. Then, a synthesis network merges the warped frames to produce the final frame. Recent advances in VFI quality have progressed with enhancements in intermediate flow predictions (Huang et al. 2022b; Kong et al. 2022; Lu et al. 2022; Zhang et al. 2023; Li et al. 2023), sparking specialized architectures to improve flow accuracy (Park et al. 2020; Park, Lee, and Kim 2021; Park, Kim, and Kim 2023). Following this direction of studies, our work aims to focus on improving intermediate flow prediction. Unlike most methods that heavily rely on reconstruction loss for end-to-end training, with optional flow distillation loss (Huang et al. 2022b; Kong et al. 2022) for stabilized training, our approach employs disentangled and direct supervision solely on flow estimation, marking an innovation in VFI research.
@@ 2.2 Perception-oriented Restoration
Conventional restoration methods in computer vision, including VFI, focused on minimizing $L_{1}$ or $L_{2}$ distances, often resulting in blurry images (Ledig et al. 2017) due to prioritizing of pixel accuracy over human visual perception. Recent studies have shifted towards deep feature spaces for reconstruction loss (Johnson, Alahi, and Fei-Fei 2016) and evaluation metrics (Zhang et al. 2018; Ding et al. 2020), demonstrating that these align better with human judgment. These approaches emphasize perceptual similarities over traditional metrics like PSNR, signaling a move towards more visually appealing, photo-realistic image synthesis.
Ever since the pioneering work of SRGAN (Ledig et al. 2017), generative models have been actively used to enhance visual quality in restoration tasks (Saharia et al. 2022b; Menon et al. 2020). The adoption of generative models has also been explored in VFI (Voleti, Jolicoeur-Martineau, and Pal 2022; Danier, Zhang, and Bull 2024; Wu et al. 2024). LDMVFI (Danier, Zhang, and Bull 2024), closely related to our work, use latent diffusion models (Rombach et al. 2022) to enhance perceptual quality. Our method aligns with such innovations but uniquely focuses on generating optical flow maps, differing from prior generative approaches that directly model the pixel space.
@@ 2.3 Diffusion Models
Diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020) are popular generative models that consist of forward and reverse process. Initially, the forward process incrementally adds noise to the data $\textbf{x}_{0}$ over $T$ steps via a predefined Markov chain, resulting in $\textbf{x}_{T}$ that approximates a Gaussian noise. The diffused data $\textbf{x}_{t}$ is obtained through forward process:
@@
While a commonly used approach is to predict the noise as above ($\epsilon$-prediction), there are some alternatives, such as $\textbf{x}_{0}$-prediction (Ramesh et al. 2022) which predicts the data $\textbf{x}_{0}$ itself or v-prediction (Salimans and Ho 2022), beneficial for numerical stability.
Figure 2: Overview of our entire framework. The training procedure operates in two stages. Initially, we train a frame synthesis network and an optical flow model, with the latter providing pseudo-labels for the second stage. In the second stage of training, we focus on training a Motion Diffusion Model to predict bi-directional flow between frames. During inference, the Motion Diffusion Model generates flow fields given the input frame pair, which the frame synthesis model uses to generate the output.
Diffusion model synthesizes the data in an iterative manner following the backward process, resulting in high perceptual quality of image samples or video samples (Saharia et al. 2022a; Ho et al. 2022). Further, we are motivated by optical flow modeling with diffusion models in other tasks (Saxena et al. 2023; Ni et al. 2023), and aim to leverage the benefit of diffusion models for optical flow synthesis in video frame interpolation. Although an existing work employs diffusion model for video frame interpolation task (Danier, Zhang, and Bull 2024), our method synthesizes the intermediate optical flows rather than directly synthesizing the RGB frames.
@@ 3.1 Overview
In this paper, we focus on the goal of synthesizing an intermediate frame $I_{\tau}$ between consecutive frames $I_{0}$ and $I_{1}$, where $10<\tau<$ 1. Our method adopts a two-stage training scheme to disentangle the training of motion modeling and frame synthesis (Fig. 2). In the first stage, we train a frame synthesis network to synthesize an RGB frame from neighboring frames and their bi-directional flows. Then, we fine-tune the optical flow model to enhance flow quality. In the second stage, the fine-tuned flow model serves as a teacher for training the motion diffusion model. During inference, this motion diffusion model generates intermediate motion (bi-directional flow maps in specific), which the synthesis network uses to produce the final RGB frame.
3.2 Synthesis and Teacher Flow Models
We propose a synthesis network $\mathcal{S}$, designed to accurately generate an intermediate target frame using a pair of input frames and their corresponding optical flows from the target frames. Specifically, given a frame pair of $I_{0},I_{1}$, and the target intermediate frame $I_{\tau}$, we first use an optical flow model $\mathcal{F}$ to obtain the bi-directional flow from the target frame to the input frames:
@@ 3.2 Synthesis and Teacher Flow Models
where $i$ denotes the index of input frames. With the estimated flows and their corresponding frames from the target frame, we synthesize $\hat{I}_{\tau}$, which aims to recover the target frame $I_{\tau}$.
@@
Figure 3: Architecture of our motion diffusion model. The input pair frames are downsampled to an 8×8$\times$8 × smaller size and goes through a 3-level U-Net, which outputs a pair of coarse flow maps and their corresponding weight masks for upsampling. The convex upsampling layer takes the coarse flow maps and weight masks to return the full resolution flow maps.
We adopt pre-trained RAFT (Teed and Deng 2020) for optical flow model $\mathcal{F}$, and train the synthesis network $\mathcal{S}$ from scratch. We use an alternating optimization of two models $\mathcal{S}$ and $\mathcal{F}$. We first fix $\mathcal{F}$ to the pre-trained state, and train the synthesis network $\mathcal{S}$. Once the training of $\mathcal{S}$ converges, we freeze $\mathcal{S}$, and fine-tune $\mathcal{F}$. We fine-tune $\mathcal{F}$ so that it could provide better estimations as the teacher in the next stage of training. Note that the flow model $\mathcal{F}$ is not used during inference, but serves its purpose as the teacher for intermediate motion modeling described in Sec. 3.3.
@@
For optimization, we compute loss on the final synthesized output $\hat{I}_{\tau}$, with a combination of three terms. First, we use the pixel reconstruction error between the synthesized frame and the target frame: $\mathcal{L}_{1}=||I_{\tau}-\hat{I}_{\tau}||_{1}$. Following recent efforts (Danier, Zhang, and Bull 2024; Chen and Zwicker 2022), we adopt the LPIPS-based perceptual reconstruction loss $\mathcal{L}_{p}$ (Zhang et al. 2018), and also exploit the style loss $\mathcal{L}_{G}$ (Gatys, Ecker, and Bethge 2016) , as its effectiveness has been proved in a recent work (Reda et al. 2022). By combining the three loss terms, we define our perception-oriented reconstruction loss $\mathcal{L}_{s}$ for high quality synthesis:
@@
Further details on our perception-oriented reconstruction loss can be found in the Appendix.
Recurrent Synthesis
We build our synthesis network $\mathcal{S}$ to be of recurrent structure, motivated by the recent trend in video frame interpolation (Sim, Oh, and Kim 2021; Jin et al. 2023; Reda et al. 2022), due to its great efficiency. The inputs $I_{in}$ and $F_{\tau}$ are resized to various scales of lower-resolution, and by applying our synthesis module $\mathcal{G}$ recurrently from low-resolution and to higher resolutions, the output frame is synthesized in a coarse-to-fine manner. For our synthesis module $\mathcal{G}$, we use a simple 3-level hierarchy U-Net (Ronneberger, Fischer, and Brox 2015). Details on our recurrent synthesis procedure are described in the Appendix.